# Deep Neural Networks: Multi-Classification and Universal Approximation

### v1

@@ 6. Further comments and open problems
In this paper, we have demonstrated that a 2-wide deep neural network can address any classification problem with no more than $O(N)$ layers. Additionally, we have established a universal approximation theorem for $L^{p}(\Omega;\mathbb{R}_{+})$ functions, requiring a neural network width of $d+1$, thereby reducing the minimal width necessary for this task. Notably, our proofs for these theorems are fully constructive, explicitly detailing the parameters to be utilized, giving a fully geometric interpretation of the architecture employed, and providing a formal proof of each statement.
@@
On the other hand, the first step of the proof of Theorem 1.2 uses an approximation by simple functions of finite volume type on a regular set of hyperrectangles, ensuring an error of order $h$ [12]. However, more sophisticated nonlinear approximation procedures, such as those based on dyadic partitions, could achieve better convergence rates (see [12]). Nevertheless, to approximate functions using the neural network, an iterative algorithm that works on a non-uniform grid must be developed.
(5) Minimal width universal approximation theorem for more general spaces. Universal approximation theorems have been extended to Sobolev and Besov spaces, as discussed in [13, 39]. In [39], it is shown that the class $W^{s,q}$, when compactly embedded in $L^{p}$, can be approximated by neural networks. For $p\leq q$, piecewise polynomial approximations on uniform grids are applicable, which neural networks can approximate. Additionally, the neural network in [39] uses a width of $3125d+$ 31. The uniformity of the grid allows us to extend our methodology to estimate the network depth while maintaining a small width and understanding the parameters involved in the approximation. However, for $p>q$, nonlinear or adaptive methods are required. Although [39] constructs a neural network to approximate nonlinear functions on non-uniform grids, the choice of parameters lacks clear geometric intuition, and no algorithmic procedure is provided.
@@
Proof of the Proposition 4.1.
Recall that the classes under consideration are defined as follows:
@@
To prove that we can compress the $M$ classes, we proceed by induction. For the first class, i.e. $k=0$, the proof follows directly by noting that the assumptions of the statement ensure the existence of parameters $\mathcal{W}^{L_{0}},\mathcal{B}^{L_{0}}$ and $z_{0}\in\mathbb{R}^{2}$ such that $\phi(\mathcal{W}^{L_{0}},\mathcal{B}^{L_{0}},\mathcal{C}_{0})=z_{0}$.
Let us assume that the statement holds for some $10<k<M-$ 1. Thus, there exists a collection of points $\{z_{j}\}_{j=1}^{k}\subset\mathbb{R}^{2}$, $L_{1}>1$, and parameters $\mathcal{W}^{L_{1}},\mathcal{B}^{L_{1}}$ such that the input-output map satisfies
$\phi^{k}(\mathcal{W}^{L_{1}},\mathcal{B}^{L_{1}},\mathcal{C}_{j})=z_{j},\quad\text{ for every }j\in\llbracket 0,k\rrbracket$
In particular, by assumption, we have that
$\phi^{k}(\mathcal{W}^{L_{1}},\mathcal{B}^{L_{1}},\mathcal{C}_{k+1})=\hat{\mathcal{C}}_{k+1},\quad\text{where }\hat{\mathcal{C}}_{k+1}=\{z_{i}\text{ with }i\in\llbracket 1,N\rrbracket\,:\,y_{i}={k+1}\}.$
Let us prove the statement for $k+1$. Denote by $\hat{\mathcal{C}}=\phi^{k}(\mathcal{W}^{L_{1}},\mathcal{B}^{L_{1}},\mathcal{C})$. By hypothesis, there exist $\hat{z}_{k+1}\in\mathbb{R}^{2}$, $L_{2}\geq 1$, $\mathcal{W}^{L_{2}}$ and $\mathcal{B}^{L_{2}}$ such that
$\phi^{k+1}(\mathcal{W}^{L_{2}},\mathcal{B}^{L_{2}},\hat{\mathcal{C}}_{k+1})=\hat{z}_{k+1},\qquad\phi(\mathcal{W}^{L_{2}},\mathcal{B}^{L_{2}},\hat{\mathcal{C}}\setminus\hat{\mathcal{C}}_{k+1})\neq\hat{z}_{k+1},$ (A.2)
@@
$\phi^{k+1}(\mathcal{W}^{L_{2}},\mathcal{B}^{L_{2}},z^{1})\neq\phi^{k+1}(\mathcal{W}^{L_{2}},\mathcal{B}^{L_{2}},z^{2}),\quad\text{for all }z^{1},\,z^{2}\in\hat{\mathcal{C}}\setminus\hat{\mathcal{C}}_{k+1},\,z^{1}\neq z^{2}.$ (A.3)
Since the composition of input-output maps is again an input-output map of (1.1), we can compose $\phi^{k+1}$ and $\phi^{k}$. Then, by (A.2)-(A.3), the map $\hat{\phi}:=\phi^{k+1}\circ\phi^{k}$ satisfies that
$\hat{\phi}(\mathcal{W}^{L_{3}},\mathcal{B}^{L_{3}},\mathcal{C}_{j})=\hat{z}_{j}\quad\text{for every }j\in\llbracket 0,k+1\rrbracket,$ (A.4)
where $L_{3}=L_{2}+L_{1}$, $\mathcal{W}^{L_{3}}=\mathcal{W}^{L_{2}}\cup\mathcal{W}^{L_{1}}$, and $\mathcal{B}^{L_{3}}=\mathcal{B}^{L_{2}}\cup\mathcal{B}^{L_{1}}$. This concludes the induction. The result follows by taking $k=M-2$ in (A.4).
Proof of Corollary 1.2.
Let us consider the dataset $\{(x_{i},y_{i})\}_{i=1}^{N}\subset\mathbb{R}^{d}\times\{\alpha_{0},\dots,\alpha_{M-1}\}$, where $\{\alpha_{k}\}_{k=0}^{M-1}\subset\mathbb{R}$. Without loss of generality, assume that $y_{i}<y_{j}$ for every $i<j$, where $i,j\in\{0,\dots,M-1\}$. If $y_{0}\geq 0$, we conclude by applying Theorem 1.1. If $y_{0}<0$, we consider a new set of labels given by $\hat{y}_{i}=y_{i}-y_{0}$ for every $i\in\{0,\dots,M-1\}$, noting that $\hat{y}_{0}=0$. Furthermore, $\{\hat{y}_{i}\}_{i=1}^{N}\subset\{\hat{\alpha}_{k}\}_{k=0}^{M-1}\subset\mathbb{R}_{+}$. Then, according to Theorem 1.1, there exist parameters $\mathcal{W}^{L}$ and $\mathcal{B}^{L}$ such that for $L=2N+4M-1$ and $w_{\text{max}}=2$, the input-output map of (1.4) with $A_{j}^{1}=\text{Id}_{d_{j}}$ (the identity matrix in $\mathbb{R}^{d_{j}\times d_{j}}$) for all $j\in\{1,\dots,L\}$, satisfies
$\phi^{L}(\mathcal{A}^{L},\mathcal{W}^{L},\mathcal{B}^{L},x_{i})=\hat{y}_{i},\quad\text{for all }i\in\{1,\dots,N\}.$ (A.5)
@@
Similarly, for $\hat{y}_{i}$ such that $\hat{y}_{i}-y_{0}\geq 0$, we obtain $\phi^{1}(\hat{y}_{i})=(\hat{y}_{i}-y_{0})=y_{i}$. Therefore, the input-output map $\phi^{L+1}:=\phi^{1}\circ\phi^{L}$ can memorize the dataset $\{(x_{i},y_{i})\}_{i=1}^{N}$. Moreover, since the width and depth of the neural network defined by $\phi^{1}$ are $2$ and $1$, respectively, the resulting neural network defined by $\phi^{L+1}$ has a width of $2$ and a depth of $2N+4M$.
@@
Proof of Lemma B.1.
By contradiction, assume that there exist $\mathcal{H}_{1}$ and $\mathcal{H}_{2}$ in some $\mathcal{S}_{I_{n}}$ such that $F(\mathcal{H}_{1})\cap F(\mathcal{H}_{2})\neq\emptyset.$ Therefore, there exist $x_{1}\in\mathcal{H}_{1}$ and $x_{2}\in\mathcal{H}_{2}$ such that $F(x_{1})=F(x_{2})$. Since $\mathcal{H}_{1}$ and $\mathcal{H}_{2}$ are different hyperrectangles, there exists $k\in\llbracket 1,d+1\rrbracket$ such that $x^{(k)}_{1}\neq x^{(k)}_{2}$. Due to the fact that $\mathcal{H}_{1}$ and $\mathcal{H}_{2}$ belong to the same subregion, we have that
@@
Proof of Lemma 5.1.
From the definition of $p_{\eta}(x)$, we immediately have that $F(\mathcal{H}_{*})=\mathbf{0}_{d}$. For the second part of the lemma, we proceed by contradiction. Let us assume that there exist $\mathcal{H}_{1},\mathcal{H}_{2}$ such that $F(\mathcal{H}_{1})\cap F(\mathcal{H}_{2})\neq\emptyset$. If $\mathcal{H}_{1}$ and $\mathcal{H}_{2}$ belong to the same subregion, we are done due to Lemma B.1.

### v2

@@ 5. Universal approximation theorem
6. Proof of the training theorems
In this section, we provide the proof of Theorems 1.2 and 1.3 and Corollary 1.4.
Proof of Theorem 1.2.
By the definition of minimizer, we have
$\mathcal{J}_{\lambda}(\mathcal{W}_{\lambda}^{L},\mathcal{B}_{\lambda}^{L})$ $\leq\mathcal{J}_{\lambda}(\mathcal{W}^{L}_{*},\mathcal{B}^{L}_{*})=\lambda\left|\!\left|\!\left|(\mathcal{W}_{*}^{L},\mathcal{B}_{*}^{L})\right|\!\right|\!\right|^{2}_{2}+\frac{1}{N}\sum_{i=1}^{N}\operatorname{loss}\left(\phi(\mathcal{W}_{*}^{L},\mathcal{B}^{L}_{*},x_{i}),y_{i}\right)$
$=\lambda\left|\!\left|\!\left|(\mathcal{W}_{*}^{L},\mathcal{B}_{*}^{L})\right|\!\right|\!\right|^{2}_{2}+\mathcal{J}_{0}(\mathcal{W}^{L}_{*},\mathcal{B}^{L}_{*})=\lambda\left|\!\left|\!\left|(\mathcal{W}_{*}^{L},\mathcal{B}_{*}^{L})\right|\!\right|\!\right|^{2}_{2},$ (6.1)
where we have used (1.10). Now, by the definition of the functional $\mathcal{J}_{\lambda}$ and (6), we obtain
$\frac{1}{N}\sum_{i=1}^{N}\operatorname{loss}\left(\phi(\mathcal{W}^{L}_{\lambda},\mathcal{B}^{L}_{\lambda},x_{i}),y_{i}\right)\leq\mathcal{J}_{\lambda}(\mathcal{W}^{L}_{\lambda},\mathcal{B}_{\lambda}^{L})\leq\lambda\left|\!\left|\!\left|(\mathcal{W}_{*}^{L},\mathcal{B}_{*}^{L})\right|\!\right|\!\right|^{2}_{2}.$ (6.2)
Now, due to Corollary 1.1, we know that $\left|\!\left|\!\left|(\mathcal{W}_{*}^{L},\mathcal{B}_{*}^{L})\right|\!\right|\!\right|_{2}$ can be uniformly bounded with respect to $\lambda$. Therefore, taking $\lambda\rightarrow 0$ in (6.2), we conclude (1.12).
Next, by the definition of $\mathcal{J}_{\lambda}$, we obtain
$\lambda\left|\!\left|\!\left|(\mathcal{W}^{L}_{\lambda},\mathcal{B}^{L}_{\lambda})\right|\!\right|\!\right|_{2}\leq\mathcal{J}_{\lambda}(\mathcal{W}^{L}_{\lambda},\mathcal{B}^{L}_{\lambda})\leq\lambda\left|\!\left|\!\left|(\mathcal{W}_{*}^{L},\mathcal{B}_{*}^{L})\right|\!\right|\!\right|^{2}_{2}.$ (6.3)
Then, dividing (6.3) by $\lambda>0$, we deduce that the sequence $\{{(\mathcal{W}_{\lambda}^{L},\mathcal{B}_{\lambda}^{L})}\}_{\lambda>0}$ is bounded by $\left|\!\left|\!\left|(\mathcal{W}_{*}^{L},\mathcal{B}_{*}^{L})\right|\!\right|\!\right|_{2}$. Moreover, the Bolzano–Weierstrass Theorem ensures that this sequence has a subsequence that converges to some $(\mathcal{W}_{0}^{L},\mathcal{B}_{0}^{L})$. Denoting such a subsequence again by $\{{(\mathcal{W}_{\lambda}^{L},\mathcal{B}_{\lambda}^{L})}\}_{\lambda>0}$, we observe that
$\mathcal{J}_{0}(\mathcal{W}^{L}_{\lambda},\mathcal{B}^{L}_{\lambda})=\mathcal{J}_{\lambda}(\mathcal{W}^{L}_{\lambda},\mathcal{B}^{L}_{\lambda})-\lambda\left|\!\left|\!\left|(\mathcal{W}^{L}_{\lambda},\mathcal{B}^{L}_{\lambda})\right|\!\right|\!\right|_{2}\leq\mathcal{J}_{\lambda}(\mathcal{W}^{L}_{\lambda},\mathcal{B}^{L}_{\lambda}).$ (6.4)
Then, combining (6) and (6.4), and using the continuity of $\mathcal{J}_{0}$, we get
$\mathcal{J}_{0}(\mathcal{W}_{0}^{L},\mathcal{B}_{0}^{L})=\mathcal{J}_{0}\left(\lim_{\lambda\rightarrow 0}(\mathcal{W}^{L}_{\lambda},\mathcal{B}^{L}_{\lambda})\right)=\lim_{\lambda\rightarrow 0}\mathcal{J}_{0}(\mathcal{W}^{L}_{\lambda},\mathcal{B}^{L}_{\lambda})=0,$
concluding that $(\mathcal{W}_{0}^{L},\mathcal{B}_{0}^{L})$ is a minimizer of $\mathcal{J}_{0}$. Now, let us consider
$(\tilde{\mathcal{W}}_{0}^{L},\tilde{\mathcal{B}}_{0}^{L})\in\operatorname{argmin}\left\{\left|\!\left|\!\left|(\mathcal{W}^{L},\mathcal{B}^{L})\right|\!\right|\!\right|_{2}\,:\,\mathcal{J}_{0}(\mathcal{W}^{L},\mathcal{B}^{L})=0\right\},$
that is, $(\tilde{\mathcal{W}}_{0}^{L},\tilde{\mathcal{B}}_{0}^{L})$ is a parameter of minimal norm minimizing $\mathcal{J}_{0}$. In particular, since $(\tilde{\mathcal{W}}_{0}^{L},\tilde{\mathcal{B}}_{0}^{L})$ is a minimizer of $\mathcal{J}_{0}$, (1.13) implies that
$\left|\!\left|\!\left|(\mathcal{W}_{\lambda}^{L},\mathcal{B}_{\lambda}^{L})\right|\!\right|\!\right|_{2}\leq\left|\!\left|\!\left|(\tilde{\mathcal{W}}_{0}^{L},\tilde{\mathcal{B}}_{0}^{L})\right|\!\right|\!\right|_{2},\quad\text{for all }\lambda>0.$ (6.5)
Taking $\lambda\rightarrow 0$ in (6.5), we obtain
$\left|\!\left|\!\left|(\mathcal{W}_{0}^{L},\mathcal{B}_{0}^{L})\right|\!\right|\!\right|_{2}\leq\left|\!\left|\!\left|(\tilde{\mathcal{W}}_{0}^{L},\tilde{\mathcal{B}}_{0}^{L})\right|\!\right|\!\right|_{2}.$ (6.6)
If (6.6) holds with strict inequality, then we contradict the fact that $(\tilde{\mathcal{W}}_{0}^{L},\tilde{\mathcal{B}}_{0}^{L})$ is a parameter of minimal norm. Therefore, we must have
$\left|\!\left|\!\left|(\mathcal{W}_{0}^{L},\mathcal{B}_{0}^{L})\right|\!\right|\!\right|_{2}=\left|\!\left|\!\left|(\tilde{\mathcal{W}}_{0}^{L},\tilde{\mathcal{B}}_{0}^{L})\right|\!\right|\!\right|_{2},$
that is, $(\mathcal{W}_{0}^{L},\mathcal{B}_{0}^{L})$ is also a parameter of minimal norm that minimize $\mathcal{J}_{0}$. ∎
Proof of Theorem 1.3.
We begin by analyzing the deviation between $x_{i}^{j}$ and $\hat{x}^{j}_{i}$, the solutions of (1.2) and (1.15) at layer $j$. We have
$\|x^{j}_{i}-\hat{x}_{i}^{j}\|$ $\leq\|\hat{\bm{\sigma}}_{j}((\hat{W}_{j}\hat{x}_{i}^{j-1}+\hat{b}_{j})-\bm{\sigma}_{j}(W_{j}\hat{x}_{i}^{j-1}+b_{j})\|+\|\bm{\sigma}_{j}(W_{j}\hat{x}_{i}^{j-1}+b_{j})-\bm{\sigma}_{j}(W_{j}x_{i}^{j-1}+b_{j})\|$
$\leq\nu_{j}+\|W_{j}\|\|x^{j-1}_{i}-\hat{x}_{i}^{j-1}\|\leq\nu_{j}+\left|\!\left|\!\left|(\mathcal{W}_{*}^{L},\mathcal{B}_{*}^{L})\right|\!\right|\!\right|_{2}\|x^{j-1}_{i}-\hat{x}_{i}^{j-1}\|,$ (6.7)
where we have used the fact that the radii $R_{j}$ defined in (1.17) are large enough to control the deviation between the activations at each layer. We also used the Lipschitz continuity of the activation functions on compact sets.
Since $\|x^{0}_{i}-\hat{x}_{i}^{0}\|=0$, an iterative application of (6) yields
$\|\hat{\phi}(\hat{\mathcal{W}}_{\lambda}^{L},\hat{\mathcal{B}}_{\lambda}^{L},x_{i})-\phi({\mathcal{W}}^{L}_{*},{\mathcal{B}}^{L}_{*},x_{i})\|^{2}$ $=\|x^{L}_{i}-\hat{x}_{i}^{L}\|^{2}\leq\left(\sum_{j=1}^{L}\nu_{j}\left|\!\left|\!\left|(\mathcal{W}_{*}^{L},\mathcal{B}_{*}^{L})\right|\!\right|\!\right|_{2}^{L-j}\right)^{2}$
$\leq\|\nu\|_{2}^{2}\sum_{j=1}^{L}\left|\!\left|\!\left|(\mathcal{W}_{*}^{L},\mathcal{B}_{*}^{L})\right|\!\right|\!\right|_{2}^{2(L-j)}$
$=\|\nu\|_{2}^{2}\left(\frac{\left|\!\left|\!\left|(\mathcal{W}_{*}^{L},\mathcal{B}_{*}^{L})\right|\!\right|\!\right|_{2}^{2L}-1}{\left|\!\left|\!\left|(\mathcal{W}_{*}^{L},\mathcal{B}_{*}^{L})\right|\!\right|\!\right|_{2}^{2}-1}\right)=:\mathscr{E}(\nu,\mathcal{W}_{*}^{L},\mathcal{B}_{*}^{L}),$
where we used the Cauchy–Schwarz inequality and the formula for the sum of a geometric series.
Now, consider the following compact set
$\mathcal{K}_{\nu}:=\left\{(z,y)\in\mathbb{R}^{m}\times\mathbb{R}^{m}\,:\,\|z\|\leq\mathscr{E}(\nu,\mathcal{W}_{*}^{L},\mathcal{B}_{*}^{L}),\;y\in\{y_{i}\}_{i=1}^{N}\right\},$
and let us define
$\mathscr{A}_{\operatorname{loss}}(\nu,\mathcal{W}_{*}^{L},\mathcal{B}_{*}^{L}):=\max_{(z,y)\in\mathcal{K}_{\nu}}\operatorname{loss}(z+y,y).$ (6.8)
Due to the continuity and nonnegativity of $\operatorname{loss}$, it follows that $\mathscr{A}_{\operatorname{loss}}<\infty$ and $\mathscr{A}_{\operatorname{loss}}\geq 0$. Moreover, as $\|\nu\|_{2}\rightarrow 0$, we have $\mathscr{E}\rightarrow 0$, and hence the compact set $\mathcal{K}_{\nu}$ converges to
$\mathcal{K}_{0}=\left\{(z,y)\in\mathbb{R}^{m}\times\mathbb{R}^{m}\,:\,\|z\|\leq 0,\;y\in\{y_{i}\}_{i=1}^{N}\right\}.$
Consequently, we deduce that
$\mathscr{A}_{\operatorname{loss}}(0,\mathcal{W}_{*}^{L},\mathcal{B}_{*}^{L})=\max_{(z,y)\in\mathcal{K}_{0}}\operatorname{loss}(y,y)=0,$
since $\operatorname{loss}(y,y)=0$. Thus, $\mathscr{A}_{\operatorname{loss}}(\nu,\mathcal{W}_{*}^{L},\mathcal{B}_{*}^{L})\rightarrow 0$ as $\|\nu\|_{2}\rightarrow 0$.
Now, by the definition of minimizer and using (6.8), we obtain
$\hat{\mathcal{J}}_{\lambda}(\hat{\mathcal{W}}^{L}_{\lambda},\hat{\mathcal{B}}^{L}_{\lambda})$ $\leq\hat{\mathcal{J}}_{\lambda}({\mathcal{W}}^{L}_{*},{\mathcal{B}}^{L}_{*})=\lambda\left|\!\left|\!\left|({\mathcal{W}}^{L}_{*},{\mathcal{B}}^{L}_{*})\right|\!\right|\!\right|^{2}_{2}+\frac{1}{N}\sum_{i=1}^{N}\operatorname{loss}\left(\hat{\phi}({\mathcal{W}}^{L}_{*},{\mathcal{B}}^{L}_{*},x_{i}),y_{i}\right)$
$=\lambda\left|\!\left|\!\left|({\mathcal{W}}^{L}_{*},{\mathcal{B}}^{L}_{*})\right|\!\right|\!\right|^{2}_{2}+\frac{1}{N}\sum_{i=1}^{N}\operatorname{loss}\left(\hat{\phi}({\mathcal{W}}^{L}_{*},{\mathcal{B}}^{L}_{*},x_{i})-\phi({\mathcal{W}}^{L}_{*},{\mathcal{B}}^{L}_{*},x_{i})+\phi({\mathcal{W}}^{L}_{*},{\mathcal{B}}^{L}_{*},x_{i}),y_{i}\right)$
$=\lambda\left|\!\left|\!\left|({\mathcal{W}}^{L}_{*},{\mathcal{B}}^{L}_{*})\right|\!\right|\!\right|^{2}_{2}+\frac{1}{N}\sum_{i=1}^{N}\operatorname{loss}\left(\left[\hat{\phi}({\mathcal{W}}^{L}_{*},{\mathcal{B}}^{L}_{*},x_{i})-\phi({\mathcal{W}}^{L}_{*},{\mathcal{B}}^{L}_{*},x_{i})\right]+y_{i},y_{i}\right)$
$\leq\lambda\left|\!\left|\!\left|({\mathcal{W}}^{L}_{*},{\mathcal{B}}^{L}_{*})\right|\!\right|\!\right|^{2}_{2}+\frac{1}{N}\sum_{i=1}^{N}\mathscr{A}_{\operatorname{loss}}(\nu,\mathcal{W}_{*}^{L},\mathcal{B}_{*}^{L})=\lambda\left|\!\left|\!\left|({\mathcal{W}}^{L}_{*},{\mathcal{B}}^{L}_{*})\right|\!\right|\!\right|^{2}_{2}+\mathscr{A}_{\operatorname{loss}}(\nu,\mathcal{W}_{*}^{L},\mathcal{B}_{*}^{L}),$
which concludes the proof. ∎
Proof of Corollary 1.4.
Let us analyze the behavior of $\sigma_{\varepsilon}$ when $\varepsilon\rightarrow 0$. We first observe that
$\lim_{u\rightarrow 0^{+}}\operatorname{erf}\left(\frac{1}{u}\right)=1,\quad\text{and}\quad\lim_{u\rightarrow 0^{-}}\operatorname{erf}\left(\frac{1}{u}\right)=-1.$
Now, taking $u=\varepsilon\sqrt{2}/x$, we observe that since $\varepsilon>0$, we have
$\lim_{\varepsilon\rightarrow 0^{+}}\operatorname{erf}\left(\frac{x}{\varepsilon\sqrt{2}}\right)=\begin{cases}1&\text{if }x>0,\\ -1&\text{if }x<0.\end{cases}$
Consequently, we have
$\lim_{\varepsilon\rightarrow 0^{+}}\sigma_{\varepsilon}(x)=\lim_{\varepsilon\rightarrow 0^{+}}\frac{x}{2}\left(1+\operatorname{erf}\left(\frac{x}{\varepsilon\sqrt{2}}\right)\right)=\begin{cases}x&\text{if }x>0,\\ 0&\text{if }x<0,\end{cases}=\operatorname{ReLU}(x)=\sigma(x).$ (6.9)
Therefore, since (6.9) holds for every $x$, we deduce that $\sigma_{\varepsilon}\rightarrow\sigma$ uniformly in $\mathbb{R}$. The rest of the proof follows as a direct application of Theorem 1.3, by observing that when $\varepsilon\rightarrow 0$ then $\nu\rightarrow 0$ as well. ∎
@@ 7. Further comments and open problems
In this paper, we have demonstrated that a 2-wide deep neural network can address any classification problem with no more than $O(N)$ layers. Additionally, we have established a universal approximation theorem for $L^{p}(\Omega;\mathbb{R}_{+})$ functions, requiring a neural network width $d+1$. Notably, our proofs are fully constructive, explicitly detailing the parameters to be utilized, giving a fully geometric interpretation of the architecture employed, and providing a formal proof of each statement.
These explicit constructions yield bounds on the parameters and provide a priori estimates for minimizers of standard regularized training loss functionals in supervised learning. As the regularization parameter vanishes, the trained networks converge to exact classifiers.
@@
On the other hand, the first step of the proof of Theorem 1.4 uses an approximation by simple functions of finite volume type on a regular set of hyperrectangles, ensuring an error of order $h$ [11]. However, more sophisticated nonlinear approximation procedures, such as those based on dyadic partitions, could achieve better convergence rates (see [11]). Nevertheless, to approximate functions using the neural network, it would be necessary to develop an iterative algorithm operating over a non-uniform grid.
(5) Minimal width universal approximation theorem for more general spaces. Universal approximation theorems have been extended to Sobolev and Besov spaces, as discussed in [12, 37]. In [37], it is shown that the class $W^{s,q}$, when compactly embedded in $L^{p}$, can be approximated by neural networks. For $p\leq q$, piecewise polynomial approximations on uniform grids are applicable, which neural networks can approximate. Additionally, the neural network in [37] uses a width $3125d+$ 31. The uniformity of the grid allows us to extend our methodology to estimate the network depth while maintaining a small width and understanding the parameters involved in the approximation. However, for $p>q$, nonlinear or adaptive methods are required. Although [37] constructs a neural network to approximate nonlinear functions on non-uniform grids, the choice of parameters lacks clear geometric intuition, and no algorithmic procedure is provided.
@@
Proof.
(Proof of Proposition 4.1) Recall that the classes under consideration are defined as follows:
@@
To prove that we can compress the $M$ classes, we proceed by induction. For the first class, i.e. $k=0$, the proof follows directly by noting that the assumptions of the statement ensure the existence of parameters $\mathcal{W}^{L_{0}},\mathcal{B}^{L_{0}}$ and $\tilde{z}_{0}\in\mathbb{R}^{2}$ such that $\phi(\mathcal{W}^{L_{0}},\mathcal{B}^{L_{0}},\mathcal{C}_{0})=\tilde{z}_{0}$.
Let us assume that the statement holds for some $10<k<M-$ 1. Thus, there exists a collection of points $\{z_{j}\}_{j=0}^{k}\subset\mathbb{R}^{2}$, $\tilde{L}_{1}>1$, and parameters $\mathcal{W}^{\tilde{L}_{1}},\mathcal{B}^{\tilde{L}_{1}}$ such that the input-output map satisfies
$\phi_{k}(\mathcal{W}^{\tilde{L}_{1}},\mathcal{B}^{\tilde{L}_{1}},\mathcal{C}_{j})=z_{j},\quad\text{ for every }j\in\llbracket 0,k\rrbracket.$
In particular, we have that
$\phi_{k}(\mathcal{W}^{\tilde{L}_{1}},\mathcal{B}^{\tilde{L}_{1}},\mathcal{C}_{k+1})=\hat{\mathcal{C}}_{k+1},\quad\text{where }\hat{\mathcal{C}}_{k+1}=\{z_{i}\text{ with }i\in\llbracket 1,N\rrbracket\,:\,y_{i}={k+1}\}.$
Let us prove the statement for $k+1$. Denote by $\hat{\mathcal{C}}=\phi^{k}(\mathcal{W}^{\tilde{L}_{1}},\mathcal{B}^{\tilde{L}_{1}},\mathcal{C})$. By hypothesis, there exist $\hat{z}_{k+1}\in\mathbb{R}^{2}$, $\tilde{L}_{2}\geq 1$, $\mathcal{W}^{\tilde{L}_{2}}$ and $\mathcal{B}^{\tilde{L}_{2}}$ such that
$\phi_{k+1}(\mathcal{W}^{\tilde{L}_{2}},\mathcal{B}^{\tilde{L}_{2}},\hat{\mathcal{C}}_{k+1})=\hat{z}_{k+1},\qquad\phi_{k+1}(\mathcal{W}^{\tilde{L}_{2}},\mathcal{B}^{\tilde{L}_{2}},\hat{\mathcal{C}}\setminus\hat{\mathcal{C}}_{k+1})\neq\hat{z}_{k+1},$ (A.2)
@@
$\phi_{k+1}(\mathcal{W}^{\tilde{L}_{2}},\mathcal{B}^{\tilde{L}_{2}},z^{1})\neq\phi_{k+1}(\mathcal{W}^{\tilde{L}_{2}},\mathcal{B}^{\tilde{L}_{2}},z^{2}),\quad\text{for all }z^{1},\,z^{2}\in\hat{\mathcal{C}}\setminus\hat{\mathcal{C}}_{k+1},\,z^{1}\neq z^{2}.$ (A.3)
Since the composition of input-output maps is again an input-output map of (1.2), we can compose $\phi_{k+1}$ and $\phi_{k}$. Then, by (A.2)-(A.3), the map $\hat{\phi}:=\phi_{k+1}\circ\phi_{k}$ satisfies that
$\hat{\phi}(\mathcal{W}^{\tilde{L}_{3}},\mathcal{B}^{\tilde{L}_{3}},\mathcal{C}_{j})=\hat{z}_{j}\quad\text{for every }j\in\llbracket 0,k+1\rrbracket,$ (A.4)
where $\tilde{L}_{3}=\tilde{L}_{2}+\tilde{L}_{1}$, $\mathcal{W}^{\tilde{L}_{3}}=\mathcal{W}^{\tilde{L}_{2}}\cup\mathcal{W}^{\tilde{L}_{1}}$, and $\mathcal{B}^{\tilde{L}_{3}}=\mathcal{B}^{\tilde{L}_{2}}\cup\mathcal{B}^{\tilde{L}_{1}}$. This concludes the induction. The result follows by taking $k=M-2$ in (A.4). ∎
Proof.
(Proof of Corollary 1.3) Let us consider the dataset $\{(x_{i},y_{i})\}_{i=1}^{N}\subset\mathbb{R}^{d}\times\{\alpha_{0},\dots,\alpha_{M-1}\}$, where $\{\alpha_{k}\}_{k=0}^{M-1}\subset\mathbb{R}$. Without loss of generality, assume that $y_{i}<y_{j}$ for every $i<j$, where $i,j\in\{0,\dots,M-1\}$. If $y_{0}\geq 0$, we conclude by applying Theorem 1.1. If $y_{0}<0$, we consider a new set of labels given by $\hat{y}_{i}=y_{i}-y_{0}$ for every $i\in\{0,\dots,M-1\}$, noting that $\hat{y}_{0}=0$. Furthermore, $\{\hat{y}_{i}\}_{i=1}^{N}\subset\{\hat{\alpha}_{k}\}_{k=0}^{M-1}\subset\mathbb{R}_{+}$. Then, according to Theorem 1.1, there exist parameters $\mathcal{W}^{L}$ and $\mathcal{B}^{L}$ such that for $L=2N+4M-1$ and $w_{\text{max}}=2$, the input-output map of (1.7) with $A_{j}^{1}=\text{Id}_{d_{j}}$ (the identity matrix in $\mathbb{R}^{d_{j}\times d_{j}}$) for all $j\in\{1,\dots,L\}$, satisfies
$\phi^{L}(x_{i})=\phi(\mathcal{A}^{L},\mathcal{W}^{L},\mathcal{B}^{L},x_{i})=\hat{y}_{i},\quad\text{for all }i\in\{1,\dots,N\}.$ (A.5)
@@
Similarly, for $\hat{y}_{i}$ such that $\hat{y}_{i}-y_{0}\geq 0$, we obtain $\phi^{1}(\hat{y}_{i})=(\hat{y}_{i}-y_{0})=y_{i}$. Therefore, the input-output map $\phi^{L+1}:=\phi^{1}\circ\phi^{L}$ can memorize the dataset $\{(x_{i},y_{i})\}_{i=1}^{N}$. Moreover, since the width and depth of the neural network defined by $\phi^{1}$ are $2$ and $1$, respectively, the resulting neural network defined by $\phi^{L+1}$ has a width $2$ and a depth $2N+4M$.
@@
Proof.
(Proof of Lemma B.1) By contradiction, assume that there exist $\mathcal{H}_{1}$ and $\mathcal{H}_{2}$ in some $\mathcal{S}_{I_{n}}$ such that $F(\mathcal{H}_{1})\cap F(\mathcal{H}_{2})\neq\emptyset.$ Therefore, there exist $x_{1}\in\mathcal{H}_{1}$ and $x_{2}\in\mathcal{H}_{2}$ such that $F(x_{1})=F(x_{2})$. Since $\mathcal{H}_{1}$ and $\mathcal{H}_{2}$ are different hyperrectangles, there exists $k\in\llbracket 1,d+1\rrbracket$ such that $x^{(k)}_{1}\neq x^{(k)}_{2}$. Due to the fact that $\mathcal{H}_{1}$ and $\mathcal{H}_{2}$ belong to the same subregion, we have that
@@
Proof.
(Proof of Lemma 5.1) From the definition of $p_{\eta}(x)$, we immediately have that $F(\mathcal{H}_{*})=\mathbf{0}_{d}$. For the second part of the lemma, we proceed by contradiction. Let us assume that there exist $\mathcal{H}_{1},\mathcal{H}_{2}$ such that $F(\mathcal{H}_{1})\cap F(\mathcal{H}_{2})\neq\emptyset$. If $\mathcal{H}_{1}$ and $\mathcal{H}_{2}$ belong to the same subregion, we are done due to Lemma B.1.